OpenStudy (anonymous):

Which of the following expressions is equivalent to m/n^7? a.-mn^7 b.mn-7 c.-m/n^7 d.-m/n^-7

OpenStudy (phi):

first idea: exponents do NOT change the sign of a number

OpenStudy (jojo4eva):

The answer is b right @phi

OpenStudy (jojo4eva):

I believe the answer is B

OpenStudy (phi):

Yes, it is B the rule is \[ \frac{1} {x^{-a} } = x^a\] and \[ x^{-a} = \frac{1}{x^a} \] If you "flip" an expression, you change the sign of the exponent.

OpenStudy (jojo4eva):

yerp @dom30011165 so phi explained for you lol i just gave u the answer